### v1

# Comprehensive Evaluation on Lexical Normalization: Boundary-Aware Approaches for Unsegmented Languages

Abstract
Lexical normalization research has sought to tackle the challenge of processing informal expressions in user-generated text, yet the absence of comprehensive evaluations leaves it unclear which methods excel across multiple perspectives. Focusing on unsegmented languages, we make three key contributions: (1) creating a large-scale, multi-domain Japanese normalization dataset, (2) developing normalization methods based on state-of-the-art pre-trained models, and (3) conducting experiments across multiple evaluation perspectives. Our experiments show that both encoder-only and decoder-only approaches achieve promising results in both accuracy and efficiency.
Comprehensive Evaluation on Lexical Normalization:
Boundary-Aware Approaches for Unsegmented Languages
Shohei Higashiyama  and Masao Utiyama National Institute of Information and Communications Technology, Kyoto, Japan {shohei.higashiyama,mutiyama}@nict.go.jp
1 Introduction
User-generated text (UGT) is invaluable textual content produced by users’ activities on web platforms such as review sites and social media. UGT’s informality—its frequent use of colloquial expressions—poses substantial challenges to accurate analysis in natural language processing applications. Consequently, researchers have explored lexical normalization (LN), the task of converting non-standard word forms into standard ones. LN has been actively studied particularly for space-delimited languages such as European languages Baldwin et al. (2015); van der Goot et al. (2021). In this study, we investigate LN for unsegmented languages, focusing primarily on Japanese.
Major issues in existing LN research are summarized as a lack of comprehensive evaluation, leaving unclear which methods excel under different evaluation aspects. Specifically, (i) comparative evaluations of recent model architectures are absent, and (ii) multi-perspective analyses—examining required training data size, inference cost, and domain-specific accuracy across diverse domains—have not been conducted. These issues are common to LN research but are particularly severe for underexplored unsegmented languages. This study addresses these gaps through three key contributions: (1) dataset construction, (2) method development based on cutting-edge pre-trained models, and (3) comprehensive experiments.
First, we introduce the Japanese Multi-Domain Lexical Normalization Dataset (JMLN), a large collection of 21,402 sentences drawn from a variety of UGT sources. JMLN’s size exceeds existing Japanese LN datasets Higashiyama et al. (2021b); Kondo et al. (2025), and its domain diversity surpasses that of any current LN datasets. These properties enable both the development of methods suitable for Japanese and multi-perspective evaluation.
Second, we develop LN methods based on three modern Transformer Vaswani et al. (2017) architectures—encoder-only, encoder-decoder, and decoder-only—including a novel encoder-based infilling approach, as well as variants of generative approaches. While these boundary-aware methods are tailored for unsegmented languages, they remain broadly applicable.
Third, we evaluate these methods on JMLN and an existing Thai dataset. Multi-perspective experiments on JMLN yield in-depth insights into methods’ characteristics and trade-offs, while experiments on the Thai dataset further validate cross-lingual applicability and generalizability.
Our evaluation reveals three main findings. First, compact encoder-only models deliver the highest inference throughput, and decoder-only models excel in normalization recall—while both yield high normalization precision. Second, training models on 4k–8k sentences yields reasonable precision of around 0.7, and cutting-edge decoder-only models deliver superior recall even with fewer instances. Third, domains rich with unknown informal words exhibit low performance, especially the typo-correction domain. [^1] Our dataset and code will be available at {URL}.
2 Related Work
Text normalization has been studied for some representative purposes: mapping dialectal variants to standard language Kuparinen et al. (2023), modernizing historical writings Bollmann (2019), and verbalizing semiotic expressions for text-to-speech (TTS) synthesis Zhang et al. (2019). Specifically, lexical normalization (LN) refers to the normalization task of converting UGT at the lexical level.
To date, research on UGT normalization can be broadly divided into three categories based on the dominant methodologies of each period: rule-based and statistical methods Aw et al. (2006); Choudhury et al. (2007); Han and Baldwin (2011), pre-Transformer neural methods Chrupała (2014); Ikeda et al. (2016); Lusetti et al. (2018); Lourentzou et al. (2019), and Transformer-based methods Muller et al. (2019); Samuel and Straka (2021); Bucur et al. (2021); Bikaun et al. (2024).
To achieve robust word segmentation (WS) for UGT, previous studies have often addressed LN jointly with WS, as exemplified by research on Japanese Sasano et al. (2013); Kaji and Kitsuregawa (2014); Saito et al. (2014, 2017); Higashiyama et al. (2021a), Chinese Wang et al. (2013); Qian et al. (2015), and Thai Haruechaiyasak and Kongthon (2013). Some recent studies have adopted Transformer masked language models (MLMs) Ueda et al. (2023); Pankam et al. (2023).
Decoder-only Transformer models have recently made remarkable advances and have been applied to text normalization for TTS Zhang et al. (2024b); Shen et al. (2024) and other sequence transduction tasks Kaneko and Okazaki (2023); Shi et al. (2024). However, these models remain underexplored in UGT normalization, resulting in a lack of comparative evaluation of state-of-the-art models for this task.
Regarding evaluation domains, many studies have focused on building short message and social media datasets for European Choudhury et al. (2007); Han and Baldwin (2011); Baldwin et al. (2015); Plank et al. (2020); van der Goot et al. (2021) and Asian languages Kaji and Kitsuregawa (2014); Limkonchotiwat et al. (2021); Nguyen et al. (2024); Kondo et al. (2025), which has led to extensive model development and evaluation in these domains. Some studies have focused on other domains Higashiyama et al. (2021b); Bikaun et al. (2024). However, cross-domain evaluation research covering three or more domains remains scarce.
3 Japanese Dataset Construction
For our primary target language, Japanese, existing datasets Kaji and Kitsuregawa (2014); Osaki et al. (2017); Higashiyama et al. (2021b); Kondo et al. (2025) have limited domain diversity and size—covering one or two domains with approximately 1,000 to 6,000 sentences/posts. In this study, we have constructed JMLN, as a large-scale dataset sourced from a variety of UGT.
Data Sources and Size
We sampled original texts from various sources: Q&A site, blog site, review site, recipe site, video site, online forum, and social media platform, as well as Wikipedia edit history and conversation transcriptions (details in Appendix A.2). The constructed dataset includes 21,402 sentences with 8,885 normalization instances, i.e., non-standard and standard form pairs (details in Appendix A.1). The large data size and domain diversity of our dataset are advantages over existing Japanese datasets, enabling multi-perspective evaluations, as shown in §6.
Basic Designs
We followed Higashiyama et al. (2021b)’s annotation criteria; annotation information includes word boundaries based on the short unit word criterion Maekawa et al. (2014) and word attributes such as part-of-speech, lemma, predefined word categories (details in Appendix A.3), and standard forms of non-standard words. [^2] Non-standard forms are those with distinctive orthographic features whose frequency in the reference corpus falls below a threshold, whereas standard forms are those whose frequency exceeds a threshold. See details in Appendix A.4.
Annotation Process
As data preparation, the first author extracted sentences with a reasonable length (10–300 characters) from each original 14 datasets and divided the sentences into two candidate sets: a random (Rand) set and a manually curated (Cur) set. Then, four experienced annotators, including an annotator manager, at a data annotation company performed the annotation process as follows.
$\bullet$
Sentence selection: From the Rand-set of each original dataset, annotators sequentially selected sentences unless they contained ethically problematic contents or were unclear in meaning. From each Cur-set, annotators intentionally selected sentences containing UGT-specific category words (Appendix A.3).
$\bullet$
Word information annotation: Annotators annotated the selected sentences with word information by modifying auto-analyzed results by a morphological analyzer, MeCab Kudo et al. (2004) with UniDic Den (2009).333cwj-3.1.0 (https://clrd.ninjal.ac.jp/unidic/)
$\bullet$
Standard form annotation: We adopted the separated steps for standard form (SForm) annotation—assigning an SForm ID for each non-standard word and associating a set of valid standard forms with each SForm ID, where the later step for all sentences was performed by the annotation manager.
Inter-Annotator Agreement
During the annotation process, manager A selected a total of 240 sentences from Cur-sets, and then two annotators—B and either C or D—independently annotated those sentences with boundaries and attributes. Inter-annotator agreement (IAA) for non-standard word recognition on these sentences, as measured by F1 score, was 0.836 (See Appendix A.6). [^4] Plank et al. (2020) and van der Goot et al. (2020) used Cohen’s kappa to evaluate IAA in informal word classification, based on whether each given word is normalized or not. This metric is not directly applicable to unsegmented languages. The datasets’ high annotation consistency and large size suggest its usefulness, and this is further demonstrated by the experiments in §6, where the evaluated models achieve high normalization accuracy.
4 Task Definition
Following previous studies Sasano et al. (2013); Baldwin et al. (2015), we define LN as a task of boundary-aware span extraction and conversion, in which a system not only generates a normalized text but also identifies the original spans of each informal words (or phrases). This task can be performed on any LN dataset, provided that informal-to-formal alignments are annotated. [^5] Otherwise, a non-trivial alignment processing is needed. We provide an alignment examples in Appendix A.5. Unlike the text-to-text conversion task Ikeda et al. (2016), the boundary-aware task enables fine-grained evaluation at the normalization span level and offers better interpretability of system outputs. An example input-output pair for our task is shown in Figure 1.
Formally, an LN system takes as input a source sentence, namely, a sequence of $n$ character (or subword) tokens $\bm{x}=\bm{x}_{0:n}=[x_{0},\dots,x_{n-1}]$, and is required to predict the set of non-standard word spans and their standard forms $S=\{(b,e,s)\}$. Here, $(b,e)$ ($0\!\leq\!b\!\leq\!e\!\leq\!n$) indicates a span of an non-standard word $x_{b:e}$ with length $e-b$ in the source sentence, and $s$ indicates its standard form. Each standard form $s$ is a string with length $\geq 0$,
where length $=0$ indicates that the non-standard word should be deleted in the normalized sentence. When $b=e$, a zero-length span indicates some token(s) should be inserted into the position $b$.
Figure 1: Flow of our detect&infill approach for an input text “{CJK}UTF8ipxmついったみてる,” which means “(I’m) looking at Twitter.” “M” and “S” represent the MASK and SEP token, respectively. The original characters “{CJK}UTF8ipxmついった” follow the SEP token, but are omitted in the Figure.
5 Methods
We present boundary-aware LN methods based on three Transformer architectures: encoder-only, encoder-decoder, and decoder-only. Our encoder-based infilling approach is a novel method for unsegmented languages, and comparing multiple approaches across different architectures offers a valuable, novel evaluation.
5.1 Infilling Approach
Among encoder-based methods, including text editing Ueda et al. (2023) and MLM infilling Muller et al. (2019), the latter directly leverage the capabilities of pretrained MLMs to insert any token from the vocabulary. As a representative study, Muller et al. (2019) proposed a two-step approach for space-delimited text, which predicts the infilling lengths and infilling tokens from subword-tokenized text. While we follow the two-step detect-and-infill framework, we propose a solution tailored to unsegmented languages. Our method jointly predicts word boundaries and infilling lengths from the input character sequence, thereby identifying non-standard word spans and their corresponding normalized spans, without the explicit alignment step Muller et al. (2019).
[…]
In the infilling step, the encoder takes as input the source text $\bm{x}^{\prime}$, in which tokens in non-standard word spans are replaced by the MASK tokens, and the MLM head predicts appropriate tokens for the masked positions from the hidden representations $\bm{h}^{\prime}$. We apply input extension to the masked source text by concatenating the original characters of the specified non-standard tokens, with SEP tokens inserted between them, similarly to existing sequence transduction methods Qiang et al. (2021); He et al. (2023).
We refer to the above method as the Full-Seg approach. [^7] We report preliminary experiments on additional variants in the Appendix B During training, the model is optimized using the sum of cross-entropy losses over multiple subtasks from both steps.
5.2 Generative Approaches
Encoder-decoder models have been extensively used in normalization research for text-to-text conversion, which we refer to the plain full-text (Plain) approach. To eliminate the informal-to-formal alignment step required for this approach, Bikaun et al. (2024) generates outputs in which non-standard words and their corresponding normalized forms are each surrounded by distinct special tokens. We introduce two generative approaches for encoder–decoder and decoder-only models, one of which—Struct—is essentially equivalent to Bikaun et al. (2024)’s method. Because few studies compare multiple generative approaches to LN—and none include decoder-only models—it is valuable to evaluate and contrast these methods across both architectures.
Approach Target Text
Plain {CJK}UTF8ipxmツイッターみてる
Struct {CJK}UTF8ipxm[[ついった> >ツイッター]]みてる
Span {CJK}UTF8ipxmついった> >ツイッター> > 0
Table 1: Expected output text of each generative approach for an input text “{CJK}UTF8ipxmついったみてる.”
As shown in Table 1, the structured full-text (Struct) approach generates a full normalized text with specifying the substrings before and after normalization and their spans, using symbols “[[,” “> >,” and “]].” The other normalization span-only (Span) approach generates not full-text but only substrings before and after normalization using a symbol “> >.” The number ($\geq 0$) succeeding the normalized substring represents how many times the same original substring occurred before that of interest in the original text, which is used for specifying the exact span of the original substring. A symbol “||” is used to separate multiple normalization instances within the input text. Only “NONE” is output when no normalization is necessary.
Encoder-Decoder
The model is trained to generate the target text from each input source text via the standard sequence-to-sequence training.
Decoder-only
[…]
Lang Dataset Set #Sent #Norm
ja JMLN train 13,196 5,879
dev 1,880 791
test-C 3,786 1,705
test-R 2,540 510
th VISTEC-2021 train 40,000 130,790
test 10,000 32,819
Table 2: Dataset statistics: The number of sentences (#Sent) and normalization instances (#Norm).
6 Experiments
We set the following experimental questions (EQs):
  1. 1.
    Across different model architectures, backbone models, and normalization approaches, which methods excel in normalization accuracy (precision and recall), and which methods are efficient in terms of inference cost?
  2. 2.
    How many training instances are required to achieve reasonable performance, and does this requirement vary by methods?
  3. 3.
    What domains and other instance characteristics are particularly challenging to normalize?
To address these EQs, we evaluated various LN methods using our Japanese dataset and an existing Thai dataset. Specifically, experiments in §7.17.3, §7.4, and §7.57.6 correspond to EQ1, EQ2, and EQ3, respectively.
As evaluation metrics, we use precision and recall at the normalization-span level for individual non-standard words, as well as the F0.5 score, which emphasizes precision over recall (see details in Appendix C.1).
Backbone Size Full-Seg Approach Struct Approach Span Approach
P R F0.5 P R F0.5 P R F0.5
E BERT-base 91M 0.713 0.529 0.667
E RoBERTa-base 100M 0.718 0.523 0.669
E DeBERTa-base 100M 0.729 0.506 0.670
E BERT-large 310M 0.762 0.570 0.714
E RoBERTa-large 320M 0.755 0.550 0.702
E DeBERTa-large 330M 0.750 0.568 0.705
S T5-base 250M 0.701 0.459 0.634 0.689 0.508 0.643
S mT5-base 580M 0.606 0.406 0.551 0.655 0.427 0.592
S T5-large 780M 0.728 0.509 0.670 0.704 0.525 0.659
S mT5-large 1.2B 0.600 0.421 0.553 0.718 0.467 0.648
D Llama-3.2-1B 1.2B 0.654 0.489 0.612 0.626 0.480 0.590
D Sarashina2.2-1B 1.4B 0.722 0.612 0.697 0.761 0.585 0.717
D Qwen2.5-1.5B 1.5B 0.623 0.459 0.580 0.338 0.516 0.363
D TinySwallow-1.5B 1.5B 0.605 0.556 0.593 0.667 0.569 0.645
D Qwen2.5-3B 3.1B 0.598 0.537 0.583 0.376 0.537 0.400
D Llama-3.2-3B 3.2B 0.680 0.522 0.641 0.666 0.530 0.633
D Sarashina2.2-3B 3.4B 0.774 0.668 0.751 0.781 0.660 0.754
D Sarashina2-7B 7.3B 0.743 0.649 0.722 0.744 0.657 0.724
D Qwen2.5-7B 7.6B 0.717 0.558 0.678 0.379 0.589 0.408
D Llama-3.1-8B 8.0B 0.738 0.538 0.687 0.719 0.549 0.677
[…]
Table 3: JMLN test results of Japanese LN models (E: encoder, S: seq2seq, D: decoder). $\hookrightarrow$” indicates the continual pre-trained model derived from the base model listed in the previous row. The best score within each size group is shown in bold. For each backbone model, the better of the Struct and Span approaches is underlined. Scores where the Span approach shows a +5%, +10%, or -10% increase/decrease compared to the Struct approach are highlighted with $\blacksquare$ blue, $\blacksquare$ light blue, and $\blacksquare$ pink backgrounds, respectively.
6.1 Datasets
Table 2 shows the statistics of two experimental datasets: JMNL and Thai VISTEC-2021 Limkonchotiwat et al. (2021). [^8] https://github.com/mrpeerat/OSKut For JMLN, we divided the Cur-set for each domain (01–14) into train, dev, and test-C sets, and merged the all train, dev, and test-C sets into unified train, dev, and test-C sets across all domains, respectively. [^9] We created the unified Rand-set across all registers similarly, but omit experimental results on this set. For VISTEC-2021, we followed the provided training/test split and regarded randomly-sampled 5% sentences in the training set as a dev set.
6.2 Models
As the backbone of LN systems, we used the following Japanese or multilingual pre-trained models in Japanese experiments: BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), and DeBERTa He et al. (2021) as character-level encoder-only models, T5 Raffel et al. (2020) and mT5 Xue et al. (2021) as encoder-decoder models, and Llama 3.1/3.2 Grattafiori et al. (2024), Qwen2.5 Qwen Team (2025), Lllama-3.1-Swallow Fujii et al. (2024), TinySwallow Shing et al. (2025) and Sarashina2/2.2 Intuitions (2024a, b) as decoder-only models. Similarly, we used the following pre-trained models in the Thai experiments: RoBERTa Yasuoka (2023) as a character-level encoder-only model, T5 and mT5 as encoder-decoder models, and Llama 3.1/3.2, Qwen2.5, Typhoon 2 Pipatanakul et al. (2024), OpenThaiGPT 1.5 Yuenyong et al. (2025), and SeaLLMs 3 Zhang et al. (2024a) as decoder-only models. Specific model instances are listed in Appendix C.2.
We fine-tuned each model, with applying LoRA Hu et al. (2022) or QLoRA Dettmers et al. (2023) to some decoder-only models, and selected the model checkpoint with the best F0.5 score on the dev set. We fine-tuned all models twice and report mean scores for two runs. The hyperparameter settings are listed in Appendix C.3.
7 Results and Analysis
7.1 Normalization Accuracy for Japanese
We evaluated LN methods with three type of architectures on the JMLN test-C set. Table 3 shows the performance of encoder-only models, and encoder-decoder and decoder-only models with both Struct and Span approaches. [^10] The Plain approach yielded performance similar to the other two approaches on the dev set, as shown in §D.3.
The observed results are as follows. (1) Among encoder-only models, the large models outperformed base models, while different backbone models showed similar performance. (2) Among generative methods, the Span approach basically achieved performance comparable to or better than the Struct approach in many cases. [^11] The only exception is the Qwen2.5 model series; the models generated many nonsensical text flagments—such as “str1> >str2> >0,” where “str1” did not appear in the original input text—until reaching the maximum output length, resulted in low precision and long inference time. (3) Within the same model series—T5, mT5, Llama-3.2, Qwen2.5, and Sarashina2.2—performance improved with increasing model size up to 8B (See Figure in Appendix D.1). However, in our preliminary experiment on the dev set, we observed no salient additional gains from the 13B–15B models. (4) Performance within groups of similarly sized models was not equivalent; certain series—specifically Sarashina2.2 series, followed by Swallow, demonstrated saliently superior performance. (5) These strong decoder-only models outperformed models with the other two architectures in recall.
In conclusion, encoder-only models demonstrated high performance despite its small size, and Sarashina2.2-3B model achieved the highest performance overall, indicating that the high capability of the backbone model was beneficial for this task.
7.2 Normalization Accuracy for Thai
On the Thai VISTEC test set, we evaluated LN methods, with only the Span approach for generative methods. Table 4 shows the results.
Backbone Size P R F0.5
E RoBERTa-base 88M 0.713 0.529 0.666
S T5-base 250M 0.645 0.618 0.640
S mT5-base 580M 0.642 0.629 0.639
S mT5-large 1.2B 0.660 0.609 0.649
D Llama-3.2 1.2B 0.628 0.628 0.628
D $\hookrightarrow$ Typhoon2 1.2B 0.644 0.634 0.642
D SeaLLMs3 1.5B 0.462 0.689 0.495
D Qwen2.5 1.5B 0.472 0.702 0.505
D Qwen2.5 3.1B 0.457 0.706 0.492
D Llama-3.2 3.2B 0.641 0.647 0.642
D $\hookrightarrow$ Typhoon2 3.2B 0.656 0.668 0.658
D SeaLLMs3 7.6B 0.465 0.705 0.499
D Qwen2.5 7.6B 0.461 0.709 0.496
D $\hookrightarrow$ ThaiGPT1.5 7.6B 0.653 0.672 0.657
D Llama-3.1 8.0B 0.653 0.659 0.655
D $\hookrightarrow$ Typhoon2 8.0B 0.661 0.678 0.664
Table 4: VISTEC test results of Thai LN models.
Similarly to the Japanese results, performance improved with increasing model size within the same model series. Additionally, continually pre-trained models focusing on Thai outperformed their base models. Overall, the small encoder-only RoBERTa-base exhibited the best precision, while all encoder-decoder and decoder-only models surpass it in recall. This introduces a precision-recall trade-off in model selection.
Model V100 H200 V100 H200
Full-Seg
BERT-large 508.8 1420.9
RoBERTa-large 561.4 1407.2
DeBERTa-large 395.0 1038.3
Struct Span
T5-large 136.0 312.3 159.0 208.4
Sarashina2.2-1B 66.5 202.1 75.4 243.1
Sarashina2.2-3B 32.6 118.1 36.0 117.4
Sarashina2-7B 17.8 73.3 19.8 83.4
Table 5: Throughput: the number of sentences processed per second, measured on a V100 and H200 GPU.
7.3 Inference Throughput
For selected models with high normalization accuracy, we measured their inference throughput using the JMLN test set, on both an NVIDIA V100 GPU with 32 GiB memory and an H200 GPU with 140 GiB memory (See detailed settings in Appendix C.6). Table 5 shows the results.
We observed: (1) Encoder-only models were the fastest, followed by T5, the smaller Sarashina model, and finally the larger Sarashina model. (2) The Span approach yielded modest gains over Struct, except for T5 on the H200. Since instruction text occupies a large proportion of total output tokens, more concise instruction prompts can improve throughput; however, it is necessary to explore prompts that preserve normalization accuracy. (3) Sarashina models exhibited substantially lower throughput on the V100 than on the H200. Their low throughput on the V100 is a critical drawback, but they run much faster on the H200. Thus, when high-spec GPUs are available, Sarashina models are viable options in accuracy-critical scenarios.
Data Register Surf-Out-Train Avg-3M DeBERTa-L T5-L Sarashina2.2-3B
F0.5 P R P R P R
01  BJ-OC Q&A site 0.59 0.595 0.654 0.416 0.572 0.369 0.717 0.578
02  BJ-OY Blog 0.48 0.713 0.720 0.543 0.741 0.551 0.802 0.662
03  RC-BLG Blog 0.37 0.775 0.801 0.691 0.793 0.664 0.803 0.728
04  RC-REV Reviews 0.43 0.790 0.868 0.731 0.790 0.593 0.808 0.741
05  RK-ICB Reviews 0.52 0.680 0.808 0.458 0.693 0.500 0.764 0.521
06  RK-TRV Reviews 0.40 0.715 0.763 0.490 0.820 0.433 0.795 0.663
07  RK-RCP Recipes 0.35 0.792 0.815 0.684 0.817 0.663 0.834 0.734
08  AM Reviews 0.32 0.798 0.862 0.728 0.776 0.655 0.832 0.748
09  NC-VID Video desc. 0.39 0.645 0.627 0.538 0.783 0.677 0.730 0.629
10  NC-PED Forum 0.27 0.795 0.850 0.747 0.666 0.492 0.808 0.772
11  TW Social media 0.60 0.633 0.613 0.511 0.600 0.464 0.763 0.677
12  JW Wiki hist. 0.69 0.312 0.528 0.140 0.197 0.095 0.509 0.280
13  NU Conv. trans. 0.17 0.758 0.845 0.745 0.683 0.596 0.800 0.745
14  SK Conv. trans. 0.43 0.718 0.771 0.550 0.728 0.457 0.830 0.664
All 0.44 0.721 0.750 0.568 0.754 0.562 0.781 0.660
Table 6: JMLN test results of representative three models for each domain. “Surf-Out-Train” indicates the Surf-Outside-Train rate for each domain test set. Avg-3M indicates the average of F0.5 scores of the three models. (Surf-Outiside-Train rate values above 0.5 and Avg-3M values below 0.7 are highlighted by underlining.)
7.4 Investigation of Training Data Size
We generated size-$N$ training sets by sampling random $N$$\in$$\{500,1\mbox{k},2\mbox{k},4\mbox{k},8\mbox{k},12\mbox{k}\}$ sentences from the entire JMLN training set, and we then fine-tuned each of DeBERTa, T5, and Sarashina models twice for each size-$N$ training set. As shown in Figure 2, the results are as follows.
First, a general trend across all models is that precision and recall improve as the data size increases. From sizes 8k to 12k the gains are more gradual, but performance is not yet saturated. Within the evaluated range, more data yields better results; however, even a 4k to 8k-size dataset can achieve reasonable precision around 0.70 when creating large amounts of annotated data is impractical.
Second, in model-specific comparisons, precision and recall follow different patterns. Precision shows no clear differences across models. In contrast, recall is consistently highest for Sarashina2-7b, followed by Sarashina2.2-1b, and lower for DeBERTa and T5; this indicates that the Sarashina-series models generalize well in terms of coverage, even when the training data size is small.
Figure 2: JMLN test results for each training data size.
7.5 Results Across Domains
As shown in Table 6, we evaluated performance of selected models—DeBERTa-large, T5-large (Span), and Sarashina-2.2-3B (Span)—for each domain test set of JMLN.
First, to explore what makes a domain difficult, we examined an indicator: the proportion of non-standard surface tokens in the test set that are not found among the training set’s non-standard surfaces (Surf-Outside-Train rate). We then computed the Pearson correlation coefficients $r$ between the indicator and the average F0.5 scores across the three models, obtaining a strong negative correlation ($r=-0.78$). Notably, for 9 out of 10 domains with a Surf-Outside-Train rate below 0.5 had average F0.5 scores above 0.7, whereas all 4 domains with a rate above 0.5 had average scores below 0.7.
Next, model performance comparisons revealed the following. (1) In all domains, all models exhibited higher precision than recall, showing a desirable characteristic because invalid normalizations would degrade downstream task performance. (2) Across most domains, Sarashina-2.2 achieved higher recall than the other models, resulting its superior overall performance. Notably, this model achieved recall over 0.5 across typical UGT domains (01–11). (3) All models exhibited very low recall below 0.3 in domain 11—a specialized domain data originated from Tanaka et al. (2020)’s typo correction dataset. Training with the task-specific dataset would improve performance, but we leave this for future.
DeBERTa T5 Sarashina2.2
#Predictions 614.5 598.5 676.5
Norm-In-Lex rate 0.875 0.928 0.921
#TPs 469.0 449.0 535.5
Surf-In-Train rate 0.806 0.806 0.716
#FPs 145.5 149.5 141.0
Surf-In-Train rate 0.289 0.183 0.238
#FNs 322.0 342.0 255.5
Surf-In-Train rate 0.258 0.290 0.302
Table 7: Models’ prediction statistics on JLMN dev set (DeBERTa: large, T5: large, Sarashina: 3B).
7.6 Error Analysis
For the three models evaluated in §7.5, we analyzed output patterns on the JMLN dev set. Specifically, we counted predictions, true positives (TP), false positives (FP), and false negatives (FN) for each model, and measured (i) the proportion of predicted normalized forms appearing in the UniDic lexicon (Norm-In-Lex rate) and (ii) the proportion of original surface forms for TP, FP, and FN instances that matched any non-standard forms in the training set (Surf-In-Train rate). Table 7 shows the results averaged over two runs per model.
DeBERTa exhibited a notably lower Norm-In-Lex rate than the other models (0.921–0.928) and the gold standard (0.947). By manually inspecting error cases, we found that DeBERTa’s restored tokens within spans sometimes formed nonsensical words. [^12] E.g., {CJK}UTF8ipxmかど (orig. {CJK}UTF8ipxmヶド, gold {CJK}UTF8ipxmけど/{CJK}UTF8ipxmけれど, “but”) and {CJK}UTF8ipxmだじりん (orig. {CJK}UTF8ipxmだぁりん, gold {CJK}UTF8ipxmダーリン, “darling”). Both suggest that the model’s independent prediction at each MASK position makes it especially prone to such errors.
Regarding the Surf-In-Train rate, all models exhibited similar trends. For TPs, approximately 70–80% of the original surface forms were known (i.e., appeared in the training set), indicating that many correct predictions relied on the seen normalization instances. For both FPs and FNs, only approximately 20–30% were known, indicating that the majority of errors involved unseen expressions. This suggests considerable room for improving generalization in normalizing unseen cases. These results align with the findings in §7.5.
8 Conclusion
This paper presented our multi-domain Japanese LN dataset, LN methods based on three Transformer architectures for unsegmented languages, and multi-perspective experiments and analysis.
The answers to the three evaluation questions (§6) are summarized as follows. (1) Compact encoder-only models achieved high precision and offered the best inference throughput, while cutting-edge decoder-only models delivered high precision, notably high recall, and reasonable throughput on a high-spec GPU. (2) Normalization accuracy consistently increased with training data size, yet even 4k–8k training sentences yielded reasonable precision around 0.7. Sarashina-series models, in particular, achieved superior recall with fewer training sentences. (3) Domains with higher rates of unknown non-standard tokens correlated with decreased performance across models. Typo correction emerged as the most challenging category, reflecting the difficulty posed by diverse typo patterns.
In future work, we will evaluate the impact of LN on downstream tasks and explore the development of a general-purpose decoder model with robust normalization capabilities.
Limitations
Dataset Size
The experimental results in §7.1—high precision up to 0.78—indicates that our dataset is large enough to train high-accuracy models. However, results in §7.4 show no clear saturation even at the maximum training size, suggesting that additional data could further improve performance. Given the cost limits of manual annotation, a promising direction is to explore methods for generating high-quality synthetic data.
Language Coverage
We evaluated our methods only on Japanese and Thai datasets, but they are readily applicable to other unsegmented and space-delimited languages. Validation on additional languages remains future work.
Inference Throughput Settings
To ensure fair comparison, we measured throughput using a single GPU via the Hugging Face Transformers Wolf et al. (2020) library. However, throughput could be improved through multi-GPU parallelism, model quantization, or adoption of high-performance inference engine, such as vLLM Kwon et al. (2023).
Encoder-Only Architecture Variants
An state-of-the-art encoder-only model, ModernBERT Warner et al. (2024), might achieve performance on par with or exceeding the models we evaluated. Due to computational and time constraints, this remains future investigation.
Word-Level Evaluation Metrics
Our word-level normalization metrics treat any span mismatch as an error—even if the predicted normalization is semantically valid—which we observed especially in decoder-only model outputs (see Appendix D.5 for examples). Such span differences have little impact on most downstream applications, so the practical usefulness of the outputs may exceed the scores reported. To complement word-level metrics, we also provide sentence-level exact-match accuracy and the chrF score Popović (2015) in Appendix D.2 and D.2.
Ethics Statement
License of Resources
MeCab is available under GPL, LGPL, and BSD License. UniDic (“unidic-cwj-3.1.0”) is available under GPL v2.0, LGPL v2.0, and New BSD License. sacreBLEU is available under Apache License 2.0 (we used this software for preliminary experiments shown in Appendix D). The licenses for the datasets and pre-trained models are listed in Appendices C.2 and A.2 (Table 19). Our use of these resources for academic research aligns with their intended use. We will release our JMLN dataset for academic research in information science; it will include only annotation information and not the original texts.
Human Annotators
The annotation work was performed by annotators at a professional data annotation company. The payment amount to the company was based on the estimate submitted by the company. The actual annotators and the payment amount to each annotator were determined by the company. The annotation work was performed by four annotators, including an annotation manager, all of whom are native Japanese speakers. Under the contract for the annotation work, it was agreed that the intellectual property rights to the deliverables would be transferred to the authors’ institution.
Potential Risks
Appropriate normalization can facilitate NLP applications while preserving the core meaning of the original texts. However, it may diminish subtle nuances and intentions in the original text; for example, casual expressions may be rendered formal, dialectal expressions may be replaced by semantically similar standard language forms, or incorrect normalization may produce an entirely different meaning.
Appendix A JMLN Dataset
A.1 Data Statistics
The detailed dataset statistics are shown in Table 8 and Table 9.
ID Name #Sent #Word #Norm
01 BJ-OC 1,441 28,631 928
02 BJ-OY 1,785 29,092 1,445
03 RC-BLG 2,312 37,135 766
04 RC-REV 1,541 31,283 310
05 RK-ICB 1,251 21,548 251
06 RK-TRV 1,610 29,240 289
07 RK-RCP 2,479 29,834 1,104
08 AM 1,769 28,055 477
09 NC-VID 918 11,908 262
10 NC-PED 947 14,858 387
11 TW 1078 14,701 858
12 JW 540 18,817 503
13 NU 507 9,176 509
14 SK 684 11,935 286
Total 18,862 316,216 8,375
Table 8: Statistics of the JMLN Cur-sets.
ID Name #Sent #Word #Nrom
01 BJ-OC 200 3,981 33
02 BJ-OY 201 3,821 56
03 RC-BLG 200 2,903 57
04 RC-REV 200 3,872 27
05 RK-ICB 200 2,942 16
06 RK-TRV 200 3,139 18
07 RK-RCP 200 2,763 25
08 AM 200 2,932 12
09 NC-VID 150 2,418 23
10 NC-PED 182 2,672 46
11 TW 207 3,127 99
13 NU 200 2,940 82
14 SK 200 2,182 16
Total 2,540 39,692 510
Table 9: Statistics of the JMLN Rand-sets.
A.2 Data Sources and Licenses
To construct our JMLN dataset, we used following datasets and text sources as shown in Table 10:
ID Data name Source dataset Text register Year
01 BJ-OC BCCWJ: Yahoo! Chiebukuro Q&A posts/responces 2004–2005
02 BJ-OY BCCWJ: Yahoo! Blog Blog posts 2008–2009
03 RC-BLG Recruit: beauty salon blogs Blog posts 2012–2014
04 RC-REV Recruit: beauty salon reviews Reviews (beauty salon) 2012–2014
05 RK-ICB Rakuten Ichiba Reviews (EC site) 2019
06 RK-TRV Rakuten Travel Reviews (hotel site) 2017–2019
07 RK-RCP Rakuten Recipe Recipes 2017
08 AM Multilingual Amazon Reviews Corpus Reviews (EC site) 2000–2015
09 NC-VID Niconico Dataset: Video meta data Video descriptions 2018
10 NC-PED Niconico Dataset: Forum data Forum posts/replies 2008–2014
11 TW Twitter Social media posts 2020–2022
12 JW Japanese Wikipedia Typo Dataset Encyclopedia edit history –2021
13 NU Nagoya University Conversation Corpus Conv. transcriptions 2001–2003
14 SK Skype Conversation Corpus Conv. transcriptions 2012
Table 10: Data sources of JMLN. The Year column indicates the years of original text publication.
A.3 Word Category Definition
We extended the UGT-specific Japanese word categories defined by Higashiyama et al. (2021b) and assigned each word in the annotation sentences to every category that it matches. As shown in Table 11, the categories are divided into vocabulary types and variant-form types, with the latter applied non-standard word forms.
A.4 Standard/Non-Sandard From Definition
As stated in Higashiyama et al. (2021b), “there are no trivial criteria to determine which variant forms of a word are standard forms” (and non-standard forms) “because most Japanese words can be written in multiple ways.” Thus, we followed their definition on standard and non-standard forms. In brief, the definitions can be summarized as follows: standard forms are those variants whose relative frequencies in the reference corpus exceed a set threshold, while non-standard forms are identified per variant category based on falling below category-specific frequency thresholds or exhibiting distinctive orthographic features. Example non-standard forms for each category is shown in Table 11. For more detailed definitions, see §4.2 of their paper.
Example Standard forms Translation
Vocabulary type Neologisms/Slang {CJK}UTF8ipxmコピペ copy and paste
Proper names {CJK}UTF8ipxmドラクエ Dragon Quest
Onomatopoeia {CJK}UTF8ipxmキラキラ glitter
Interjections {CJK}UTF8ipxmおお oops
Dialect words {CJK}UTF8ipxmほんま truly
Foreign words {CJK}UTF8ipxmEASY easy
Ancient words* {CJK}UTF8ipxm[行く] べし should [go]
[…]
Emoticons/AA {CJK}UTF8ipxm(^−^)
Variant-form type Character type variants {CJK}UTF8ipxmカワイイ {CJK}UTF8ipxmかわいい,可愛い cute
Alternative representations {CJK}UTF8ipxm大きぃ {CJK}UTF8ipxm大きい big
Sound change variants {CJK}UTF8ipxmおいしーい {CJK}UTF8ipxmおいしい,美味しい tasty
Typographical errors {CJK}UTF8ipxmつたい {CJK}UTF8ipxmつらい,辛い tough
Table 11: Word categories extended from Higashiyama et al. (2021b). New categories are marked with “*.” “[]” indicates the context. {CJK}UTF8ipxm[行く] にゃ” is a kitten-style sentence ending and might be expressed as “[I go,] meow” or “[Goin’]nya.” {CJK}UTF8ipxmおはこんばんにちは” is a coined blend of “good morning,” “good afternoon,” and “good evening,” and might be expressed as “Good morn-noon-evening.”
Data Number F1 agreement score
Sent Word Cate VForm Word POS Lem Cate Cateb VForm VFormb
01 BJ-OC 20 327/325 30/30 19/18 99.08 98.77 98.16 83.33 83.33 86.49 86.49
02 BJ-OY 20 359/358 52/48 17/15 99.58 96.79 95.40 86.00 94.00 75.00 81.25
03 RC-BLG 32 1424/1424 133/124 45/43 99.86 99.02 98.81 81.71 91.83 63.64 84.09
04 RC-REV 41 2802/2802 129/137 22/22 100.0 99.04 98.75 96.24 96.24 100.0 100.0
05 RK-ICB 20 600/601 27/27 15/14 99.75 98.58 98.58 85.19 96.30 82.76 89.66
06 RK-TRV 25 748/748 23/29 11/14 100.0 99.20 98.93 84.62 84.62 80.00 80.00
07 RK-RCP 22 442/442 43/39 26/20 99.32 98.64 97.96 87.80 92.68 82.61 82.61
08 AM 20 675/675 85/86 32/34 100.0 99.56 95.56 91.23 99.42 90.91 96.97
09 NC-VID 20 353/358 55/50 20/10 97.33 95.08 93.67 70.48 83.81 46.67 53.33
10 NC-PED 20 443/446 63/63 24/24 98.31 97.19 94.26 77.78 85.71 62.50 66.67
Total 240 8173/8179 640/633 231/214 99.66 98.65 97.82 85.78 92.22 76.85 83.60
Table 12: Statistics on inter-annotator agreement. The “Number” columns show counts of annotated sentences (Sent), words, word categories (Cate), variant forms (VForm) for each annotator (formatted as “value1/value2”). The “F1 agreement score” columns report F1 scores for word segmentation (Seg), parts-of-speech (POS), lemmas (Lem), Cate, and VForm. Binary agreement on the presence or absence of a category/variant-form assignment is indicated by “$b$.”
A.5 Standard Form Annotation
The annotation process first identifies word boundaries and then assigns each word to (a category and) a standard form ID, as in the example below (meaning “Super tired today.”); for simplicity, the example displays the standard-form string instead of the ID. This approach makes it possible to obtain explicit mappings between non-standard words and their standard forms.
Original text {CJK}UTF8ipxm今日はスゴクツカレタ~
Segmented text {CJK}UTF8ipxm今日 |||| は |||| スゴク |||| ツカレタ~
Standard forms {CJK}UTF8ipxm   –    ||||  –  |||| すごく |||| 疲れた
However, when only sentence-level normalization is provided, aligning standard and non-standard forms is not straightforward. In our example, the contiguous span mapping from “{CJK}UTF8ipxmスゴクツカレタ~” to “{CJK}UTF8ipxmすごく疲れた” (“super tired”) is easily identified, but determining precise word-level alignments remains challenging.
A.6 Inter-Annotator Agreement
Table 12 shows the detailed statistics on inter-annotator agreement. Sentences in domains 01–07 were annotated by Annotators B and C, and those in domains 08–10 were by B and D. Notably, after we fixed annotation disagreements for these sentences through discussions, the sentences were integrated into our final dataset.
Appendix B Encoder-based Approach Variants
In addition to the Full-Seg approach, we introduce two variant approaches based on the encoder-base detect&infill method: Part-Seg and Full-Seg-POS. Experimental results on these variants are reported in D.2.
Full/Partial Word Segmentation
[…]
Word Feature Prediction
In the detection step, we can optionally employ multi-task learning for word feature prediction. Specifically, we adopt part-of-speech (POS) tag prediction for each token using an additional linear layer if POS annotation is available. We refer to this approach as Full-Seg-POS.
Appendix C Detailed Experimental Settings
C.1 Definition of Evaluation Metrics
Assume each input sentence $\bm{x}$ has a sequence of gold normalization instances $\mathcal{N}_{\bm{x}}=\{(b_{i},e_{i},S_{i})\}_{i}$, where each instance $(b_{i},e_{i},S_{i})$ consists of a span $(b_{i},e_{i})$ and a set $S_{i}=\{s_{i,k}\}_{k=1}^{K_{i}}$ of one or more standard forms for the corresponding non-standard word. A system is required to output a set of predicted normalization instances $\hat{\mathcal{N}}_{\bm{x}}=\{(\hat{b}_{j},\hat{e}_{j},\hat{s}_{j})\}_{j}$, where each $\hat{s}_{j}$ is a single predicted standard form.
We count a predicted instance as a TP if its span $(\hat{b}_{j},\hat{e}_{j})$ matches the span $(b_{i},e_{i})$ of a gold instance and its predicted form $\hat{s}_{j}$ belongs to the corresponding gold set $S_{i}$ over all test sentences. Precision $P$, recall $R$, and the $F_{0.5}$ score over the test set are then defined as follows:
$P$ $=$ $\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}},$
$R$ $=$ $\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}},$
$F_{0.5}$ $=$ $(1+0.5^{2})\frac{PR}{0.5^{2}P+R}.$
C.2 Pretrained Models
We list all pre-trained models used in our experiments in Table 19. We selected these models based on their strong performance on general benchmarks; however, for Thai encoder-only and encoder-decoder models, few alternative candidates were available.
C.3 Model Hyperparameters
The hyperparameter values used in the experiments are listed in Table 13.
We conducted hyperparameter search for some important parameters within our computational budgets. Based on F0.5 score for the JMLN dev set, we chose the best value of learning rate from the search space of {1e-5, 2e-5, 3e-5, 4e-5, 5e-5} for encoder-based models using RoBERTa-large, chose the best value of learning rate from {1e-4, 2e-4, 3e-4, 4e-4, 5e-4} and that of beam search width during inference from {1, 2, 3} for encoder-decoder and decoder-based models using T5-large and Sarashina2-7b, respectively. We also chose the best value of LoRA rank from {4, 8, 16} for decoder-only models using Sarashina2-7b.
Hyperparameter Value
Training epochs ja: 30; th: 20
Batch size ja: 16; th: 32
Learning rate 3e-5
Learning rate scheduler linear
Warmup ratio 0.1
Gradient norm clipping threshold 1.0
Optimizer AdamW
Training epochs ja: 30; th: 15
Batch size ja: 32; th: {4, 16}
Learning rate 2e-4
Learning rate scheduler constant
Warmup ratio 0.1
Gradient norm clipping threshold 1.0
Optimizer AdaFactor
Beam width for inference 2
Training epochs 10
Batch size 8
Learning rate 2e-4
Learning rate scheduler cosine
Warmup ratio 0.03
Weight decay 0.001
Gradient norm clipping threshold 0.3
Optimizer paged_adamw_32bit
bf16 True
LoRA rank 8
LoRA alpha 16
LoRA dropout 0.05
LoRA target modules all linear layers
Quantization bit {none, 4bit}
Beam width for inference 1
Table 13: Hyperparameter settings for encoder-only (top), encoder-decoder (middle), and decoder-only models (bottom). Batch sizes of 4 and 16 were used for the Thai mT5-large model and the other Thai encoder-decoder models. For models with 7B parameters or more, 4-bit quantization were applied during fine-tuning.
C.4 Prompts for Decoder-only Models
[…]
Instruction text
If no informal {lang} word forms exist in the input text, output the text as is. Otherwise, identify informal word forms and normalize them into their corresponding standard forms. Provide the full normalized text where the original word forms are replaced with the standard forms.
If no informal {lang} word forms exist in the input text, output the text as is. Otherwise, identify informal word forms and normalize them into their corresponding standard forms. Provide the full normalized text, embedding the original and normalized word forms in the format "[[before> >after]]". Ensure that the concatenated string of the text outside the brackets and the "before" parts is identical to the input text.
If no informal {lang} word forms exist in the input text, output exactly "NONE". Otherwise, identify every informal word form and normalize it into its corresponding standard form. For each occurrence, output a record in the format "before> >after> >count". Here, count is the count of how many times the identical original string has already appeared earlier in the input text. If multiple informal forms are found, output each record in the order they occur and separate them with "||".
Figure 3: Instruction prompts for decoder-only models with Plain (top), Struct (middle), and Span approaches (bottom).
C.5 Computational Budget for Fine-tuning
In our experiments, we used NVIDIA V100 GPUs with 32GiB memory, A100 GPUs with 80GiB memory, and H200 GPUs with 140GiB memory. In total, the models were fine-tuned for 42 GPU hours on V100s, 2700 GPU hours on A100s, and 141 GPU hours on H200s. Encoder-only and encoder-decoder models were fine-tuned on a single GPU, whereas decoder-only models were fine-tuned using eight GPUs.
C.6 Throughput Calculation Setting
In the experiments reported in §7.3, we used the following settings. For each model checkpoint (from one of two runs), we conducted a single warm-up inference followed by three inference passes over all 3,786 JMLN test sentences (a total of 11.1k characters), which were sorted by increasing token count according to its tokenizer. These evaluations were run at multiple batch sizes; we selected the batch size that yielded the highest throughput (shown in Table 14) and reported the mean throughput of the three runs. Models were cast to float16 on the V100 and to bfloat16 on the H100. All inference was performed using the Hugging Face Transformers Wolf et al. (2020) library.
Model V100 H200 V100 H200
Full-Seg
BERT-large 64 128
RoBERTa-large 64 128
DeBERTa-large 64 128
Struct Span
T5-large 256 256 256 256
Sarashina2.2-1B 128 256 256 256
Sarashina2.2-3B 128 128 64 128
Sarashina2-7B 64 128 32 128
Table 14: Batch size yielding the best throughput for each method.
C.7 Word Coverage Indicators
For analyses in experiments reported in §7.5 and §7.6, we introduced three word coverage-based indicators. Below, we provide further explanations on them.
The Surf-Outside-Train rate is defined as “the proportion of non-standard surface tokens in the test set that are not found among the training set’s non-standard surfaces.” Assume that the non-standard surface tokens appearing in the entire training set are [“u”, “u”, “r”, “thx”], while those appearing in the entire test set are [“u”, “r”, “cuz”]. In this case, the rate is $1/3=0.33$.
The Surf-In-Train rate is defined as “the proportion of original surface forms for examples that matched any non-standard forms in the training set.” Assume the same training set above and non-standard surface examples of interest are [“u”, “r”, “cuz”], the rate is $2/3=0.67$
The Norm-In-Lex rate is defined as “the proportion of predicted normalized forms appearing in the UniDic lexicon.” Assume the set of normalized strings produced by the model is {“you”, “are”, “becaus”}, and an external lexicon contains only “you” and “are” from that set; the rate is $2/3=0.67$.
Appendix D Detailed Experimental Results
D.1 Accuracy Across Model Series
For the experiment reported in Table37.1), we extracted results for different model size within each model series and plotted them in Figure 4. Performance improved with increasing model size across model series.
Figure 4: Plot of F0.5 scores in Table3 for each model series—T5, mT5, Llama-3.2, Qwen2.5, and Sarashina2.2. The scores for the Struct and Span approaches are shown with solid and dotted lines, respectively.
D.2 Comparison of Encoder Approaches
As a preliminary experiment, we compared variants of encoder-based methods—Part-Seg, Full-Seg, and Full-Seg-POS, introduced in Appendix B—on the JMLN dev set. We also report the F1 scores for word segmentation, POS tagging, and length prediction (Lenp: tokens in positive non-standard words; Lenn: other tokens), as well as additional normalization metrics: sentence-level exact match accuracy (S-Accp: accuracy for sentences containing at least one positive non-standard word; S-Accn: accuracy for other sentences) and the chrF score Popović (2015) implemented in sacreBLEU Post (2018). [^21] We calculated chrF scores with the default options (signature: nrefs:1|case:mixed|eff:yes|nc:6|nw:0| space:no|version:2.5.1). Table 16 shows the results.
The encoder-only models (3 backbone models×$\times$2 model size) with the Full-Seg approach obtained +0.01–0.05 F0.5 gains from the Part-Seg counterpart, indicating the importance of learning full word segmentation tasks. Adding POS tagging task showed no clear improvements.
D.3 Comparison of Generative Approaches
As a preliminary experiment, we compared the three generative approaches—Plain, Struct, and Span— using T5, Sarashina2/2.2, and Swallow models on the JMLN dev set. The Struct and Span approaches outperformed or matched the Plain approach for T5 and Sarashina2/2.2 models (e.g., -0.009 to +0.034 S-Accp points), but underperformed for Swallow (e.g., -0.076 S-Accp points).
D.4 Results for Each Category
Table 15 shows the JMLN test performance (recall) of the three models evaluated in §7.5 for each domain. Consistent with the domain-specific evaluation results in §7.5, Sarashina achieved the highest recall across all categories. All three models showed the same recall order across the four categories, and the consistently lowest recall of the fourth category, i.e., typos, once again highlights the difficulty of correcting them.
Category # DeBERTa T5 Sara2.2
Char type var. 534 0.395 0.375 0.563
Alter rep. 271 0.550 0.500 0.703
Sound change 794 0.796 0.739 0.812
Typos 178 0.087 0.076 0.219
Table 15: JMLN test recall of representative models (DeBERTa-large, T5-large, and Sarashina2.2-3b) for each category.
D.5 Prediction Examples
Table 18 shows model outputs on the JMLN dev set for the three models evaluated in §7.5; in addition to error type classification (TP, FP, and FN) based on the gold standard evaluation, the first author also assessed the validity of each output ($\checkmark$: valid, $\triangle$: questionable, ✗: invalid).
In example (a), all three methods produced semantically valid normalized strings at the phrase level; however, only DeBERTa was classified as a TP, as it correctly matched the gold span. Similarly, in examples (b) and (c), Sarashina produced plausible normalized strings, but due to mismatches with the gold annotation, the outputs were classified as FPs. In example (d), Sarashina normalized the input to a synonym (i.e., “darling” to “husband”); while the result was semantically appropriate, it falls outside the scope of the task, which requires normalization to variant surface forms of the same word. In examples (e) and (f), Sarashina produced normalized strings that were semantically unrelated to the original input. DeBERTa generated invalid, non-word outputs in examples (b) and (d), while T5 did so in example (c). Example (g) shows that all three models produced different but erroneous outputs, suggesting that normalization becomes particularly challenging when multiple non-standard words appear in sequence, making boundary and word identification more difficult.
Backbone Approach Seg POS Lenp Lenn Norm
F1 F1 F1 F1 P R F0.5 S-Accp S-Accn chrF
Leave-As-Is 0 0 1 95.29
BERT-B P-Seg 0.644 0.676 0.510 0.635 0.487 0.982 96.69
F-Seg 0.980 0.656 0.978 0.721 0.546 0.678 0.518 0.979 96.82
F-Seg-POS 0.978 0.963 0.636 0.976 0.649 0.532 0.621 0.508 0.970 96.56
RoBERTa-B P-Seg 0.630 0.662 0.515 0.626 0.488 0.976 97.04
F-Seg 0.982 0.669 0.980 0.726 0.537 0.678 0.502 0.981 97.22
F-Seg-POS 0.982 0.971 0.677 0.981 0.713 0.566 0.678 0.525 0.973 97.27
DeBERTa-B P-Seg 0.637 0.678 0.516 0.638 0.489 0.977 97.12
F-Seg 0.982 0.656 0.980 0.745 0.516 0.685 0.477 0.989 97.16
F-Seg-POS 0.983 0.972 0.680 0.982 0.716 0.566 0.680 0.528 0.974 97.26
BERT-L P-Seg 0.682 0.717 0.570 0.682 0.542 0.979 96.92
F-Seg 0.984 0.704 0.983 0.769 0.588 0.725 0.549 0.979 97.01
F-Seg-POS 0.980 0.967 0.691 0.979 0.756 0.576 0.712 0.534 0.984 97.04
RoBERTa-L P-Seg 0.671 0.706 0.557 0.670 0.515 0.980 97.27
F-Seg 0.983 0.701 0.982 0.753 0.571 0.708 0.542 0.980 97.34
F-Seg-POS 0.985 0.974 0.713 0.983 0.765 0.580 0.719 0.551 0.984 97.49
DeBERTa-L P-Seg 0.685 0.763 0.556 0.710 0.523 0.985 97.34
F-Seg 0.986 0.718 0.984 0.763 0.593 0.722 0.556 0.980 97.54
F-Seg-POS 0.985 0.975 0.727 0.983 0.788 0.592 0.739 0.548 0.984 97.50
Table 16: Detailed results of encoder-only models on the JMLN dev sets. P-Seg, F-Seg, and F-Seg-POS indicate the Part-Seg, Full-Seg, and Full-Seg-POS approaches.
Backbone Approach P R F0.5 S-Accp S-Accn chrF
Leave-As-Is 0 0 1 95.29
T5-base Plain 0.497 0.958 96.54
Struct 0.727 0.491 0.663 0.465 0.983 96.94
Span 0.707 0.527 0.662 0.509 0.974 97.01
T5-large Plain 0.535 0.950 96.44
Struct 0.769 0.558 0.714 0.531 0.984 97.31
Span 0.751 0.568 0.705 0.566 0.975 97.28
Sarashina2.2-1B Plain 0.602 0.985 97.92
Struct 0.750 0.635 0.724 0.602 0.972 97.72
Span 0.770 0.609 0.730 0.593 0.975 97.59
Sarashina2.2-3B Plain 0.656 0.980 98.17
Struct 0.792 0.682 0.766 0.651 0.972 97.94
Span 0.792 0.677 0.767 0.653 0.979 98.11
Sarashina2-7B Plain 0.646 0.982 98.08
Struct 0.766 0.670 0.745 0.658 0.971 97.99
Span 0.767 0.662 0.743 0.637 0.965 97.74
Llama-3.1-Swallow-8B Plain 0.638 0.983 98.03
Struct 0.771 0.609 0.732 0.584 0.980 97.68
Span 0.755 0.602 0.718 0.562 0.974 97.50
Table 17: Detailed results of encoder-decoder and decoder-only models on the JMLN dev sets.
[…]
      (Not sure about the effect.)
[…]
      (Can you just was daikon radish sprouts with water and eat it as is?)
[…]
T5-L FN
[…]
      (Isn’t Kintetsu Railway’s ‘Mawaryanse” pass a good deal?)
Gold
DeBERTa-L
[…]
      (My darling treated me.)
[…]
T5-L FN
[…]
      (I’m going kan-tsuri (= fishing at a managed sport) at Walton and bass fishing
      in Tone River system.)
Gold
DeBERTa-L
T5-L
[…]
      (The high is like 25{CJK}UTF8ipxm℃ or something, right?)
[…]
DeBERTa-L FN
T5-L FN
[…]
      (Ugh, Hatsune Miku is trending again.)
[…]
ID Pretrained Model Lang Exp Size Hugging Face ID License
ja th
01 BERT-base ja $\checkmark$ 91M tohoku-nlp/bert-base-japanese-char-v3 Apache 2.0
02 BERT-large ja $\checkmark$ 310M tohoku-nlp/bert-large-japanese-char-v2 Apache 2.0
03 RoBERTa-base ja $\checkmark$ 100M ku-nlp/roberta-base-japanese-char-wwm CC BY-SA 4.0
04 RoBERTa-large ja $\checkmark$ 320M ku-nlp/roberta-large-japanese-char-wwm CC BY-SA 4.0
05 RoBERTa-base th $\checkmark$ 88M KoichiYasuoka/roberta-base-thai-char Apache 2.0
06 DeBERTa-base ja $\checkmark$ 100M ku-nlp/deberta-v2-base-japanese-char-wwm CC BY-SA 4.0
07 DeBERTa-large ja $\checkmark$ 330M ku-nlp/deberta-v2-large-japanese-char-wwm CC BY-SA 4.0
08 T5-base ja $\checkmark$ 250M retrieva-jp/t5-base-long CC BY-SA 4.0
09 T5-large ja $\checkmark$ 780M retrieva-jp/t5-large-long CC BY-SA 4.0
10 T5-base th $\checkmark$ 250M kobkrit/thai-t5-base N/A
11 mT5-base M $\checkmark$ $\checkmark$ 580M google/mt5-base Apache 2.0
12 mT5-large M $\checkmark$ $\checkmark$ 1.2B google/mt5-large Apache 2.0
13 Llama-3.2-1B M $\checkmark$ $\checkmark$ 1.2B meta-llama/Llama-3.2-1B Llama 3.2
14 Llama-3.2-3B M $\checkmark$ $\checkmark$ 3.2B meta-llama/Llama-3.2-3B Llama 3.2
15 Llama-3.1-8B M $\checkmark$ $\checkmark$ 8.0B meta-llama/Llama-3.1-8B Llama 3.1
[…]
17 Qwen2.5-1.5B M $\checkmark$ $\checkmark$ 1.5B Qwen/Qwen2.5-1.5B Apache 2.0
18 Qwen2.5-3B M $\checkmark$ $\checkmark$ 3.1B Qwen/Qwen2.5-3B Qwen Research
19 Qwen2.5-7B M $\checkmark$ $\checkmark$ 7.6B Qwen/Qwen2.5-7B Apache 2.0
[…]
21 Sarashina2.2-1B ja&en $\checkmark$ 1.4B sbintuitions/sarashina2.2-1b MIT
22 Sarashina2.2-3B ja&en $\checkmark$ 3.4B sbintuitions/sarashina2.2-3b MIT
23 Sarashina2-7B ja&en $\checkmark$ 7.3B sbintuitions/sarashina2-7b MIT
[…]